compute the force of gravity between two 0.7 kg masses whose centers are 15 centimeters apart
1
Expert's answer
2018-03-12T16:37:08-0400
The velocity of the wheel
v=ωR
ω=2πn
R=d/2
So
v=2πn d/2=πdn=3.14×1.2×1=3.77 m/s